OpenStudy (anonymous):

How does 3/8(x-1/9)=1/6(x+2) convert to 3/8x-1/24+1/6x+1/3?

OpenStudy (anonymous):

it doesn't? it would be written as 3/8x-1/24=1/6x+1/3; just multiply through, 3/72=1/24;

OpenStudy (anonymous):

Hmm..I don't know how to type in fractions here. Those are fractions. Does that make more sense?

OpenStudy (anonymous):

okay so what i'm seeing is \[3/8(x-1/9)=1/6(x+2) \] ?

OpenStudy (anonymous):

Right. I'm doing a step-by-step practice that says it simplifies to 3/8x-1/24=1/6x+1/3, but I don't understand how the program got that. It doesn't tell me.

OpenStudy (anonymous):

okay, well that is correct. so look at the left side first. \[3/8(x-1/9)\] multiply the 3/8 by the first and the second term so.. \[3/8*x-(3/8)(1/9)\] when multiplying fractions you multiply the numerators together and the denominators together. so \[(3*1)/(8*9) = 3/72\] which can be simplified to \[1/24\]. so the left side now looks like \[3/8*x-1/24\] does that make sense?

OpenStudy (anonymous):

Ok, yes. Thank you.

OpenStudy (anonymous):

right, well if that works the right side of the equation is done exactly the same way

OpenStudy (anonymous):

It worked. Thank you.